ophelia111:

help pls If the area of a square is 100 sq ft, find the ratio of the area to perimeter of a square. 3:6 2:5 5:2 3:5

Luigi0210:

Well, to start off, they gave you the area as \( 100 ft^2 \). If you use the formula \(A=s^2 \), you can find the length of one side: \[\large s^2=100 \] \[\large s=\sqrt{100}=10 \] The perimeter of a square is just the four sides added: \(\large P=4s \) So plugging in 10 you get: \[\large P=4(10) =40\] To find the ratio of area to perimeter: \[\large 100:40 \] Just simplify that

ophelia111:

ok ty

Luigi0210:

No problem, good luck (: